# The cognitive triple-slit experiment

### v2

@@ 4 The cognitive triple-slit experiment
In this section, we explain how our cognitive experiment was conducted in practice. For simplicity, we only considered a one-dimensional detection screen. Note that the typical interference fringes are a consequence of the slits’ geometry, in the sense that it is their height, and the fact that they are parallel to each other, that are responsible for the fringe appearance, while it is their distance from each other, and width, that are responsible for the alternation between high and low impact zones and diffraction phenomena. For the purposes of our experiment, however, it was not necessary to investigate the vertical dimension of the slits. More precisely, our detection screen consisted of $n=8d+3$ elementary cells of unit length, where $d=5$ corresponds to the distance between the slits (the number of cells separating them) in the slit plate. The latter was placed at a distance of $L=12$ cells from the detection screen, which was formed of $n=43$ cells, with discrete positions $x=1,\dots,n$. The slits were chosen to be of unit width and placed in correspondence with cells with numbers $n=16,22,28$ (see Figures 4 and 6).
@@
More precisely, after agreeing to participate, each participant had to read an introductory text explaining the context of the experiment. Note that the introductory text, and subsequent explanatory texts, did not mention that it was a cognitive simulation of a quantum physics experiment, so as not to condition their choices on the basis of possible prior knowledge of the double-slit and possibly triple-slit experiments. For this reason, the different situations were presented to them in purely imaginative terms, as a description of an animal looking for food to eat (see below). In addition to the introductory text, each of the tasks was preceded by an explanation of the criterion by which they had to select a certain number of points (cells) on the detection screen. Basically, in each task participants had to select $7$ cells in sequence. The first selection corresponded to what was considered the best cell to answer the question asked, receiving (during analysis) a score of $7$ points. The second selected cell, which could not be the same because once a cell was selected the system disabled it and could not be selected a second time, received a score of $6$ points, and so on, until the seventh selected cell, which received only $1$ point. In other words, respondents answered the questions using a rank-ordering procedure, with scores from $7$ to $1$ assigned during analysis. This is of course not exactly what happens in a physics laboratory, where the measuring instrument records only one impact at a time. However, if each subject had been asked to select only one cell per task, the number of data collected might have been insufficient. However, the patterns highlighted in this work remain identifiable regardless of the score that is attributed to the sequential choices of the participants, as we will explain in Section 5.
@@
For the single-slit experiments, participants were asked to select cells that were the best examples of impacts caused by particles emerging from the slit in question. For the experiments with two and three slits, they were asked to select cells that were the best examples of a situation of doubt about the slit from which the particle emerged. To best convey these situations, the various questions were posed in the form of a bet. Note that the source of particles behind the slit plate was not described, as this would have complicated the understanding of the experiment and introduced considerations of the position of the source relative to the three slits. For the participants, each slit played the role of a potential source, with the constraint that only one at a time could emit a particle. More precisely, following some preliminary questions, participants had to read the following text (see Figure 4): This test evaluates how you interpret situations involving varying degrees of uncertainty. You will be presented with different situations and for each one you will be asked some questions. Although a question may seem similar to the previous one, you must always remain focused and answer as if it were the first time. The following description ensued: There is a large rectangular area. On one side there is a row of $43$ identical portions of food. On the opposite side is a wall with one or more openings. A hungry animal emerges from an opening and moves towards the line of food. Once at the line, the animal eats the portion of food in front of it and goes away. You have no information about the nature of the animal. You don’t know if the animal can sense the food, and if it can, you don’t know which sense it uses (sight, smell, etc.). This means that its path could even be exploratory or erratic.
@@
Participants had then to execute their first task, answering the following question: In this situation, the animal can only come out of one opening. You have bet a lot of money that you can guess which portion of food the animal will eat. What is your choice? To select a portion of food, click on it. Next, they were asked to provide their best second choice, and so on, until the seventh. Then, the same situation was presented to them two more times, changing the place of the opening. Following these three single-opening (single-slit) experiments, participants had to move to their fourth task, which was different from the previous ones. The instruction for the fourth task was as follows: Many things have changed in this new situation: there are now two openings from which the animals can come out; now is you who decide which portions of food were eaten by the animals; each time a portion of food has been eaten, and the animal has left, another person, your opponent, will observe the situation and try to guess which opening the animal came from, taking into account which portion of food has been eaten; you have bet a lot of money that your opponent will not guess correctly; which portion of food should you choose for the animal to eat, so that your opponent cannot guess which opening it came from, and you win a lot of money? To select a portion of food, click on it. Note: Your opponent understands that there is uncertainty as to how the animal senses and reaches the food. In other words, in his/her reasoning, your opponent takes into account the fact that the animal may follow exploratory or irregular paths. Again, participants had to sequentially provide their best seven choices, from best to worst. Then, the same situation was presented to them two more times, changing the place of the two openings. Following these three double-opening (double-slit) experiments, participants had to move to their seventh and last task. The instruction for it was as follows: This is the last variation. There are now three openings for the animals to come out of. Which portions of food should you choose for the animals to eat, so that your opponents (different each time) cannot guess which opening they came from, and you win a lot of money? As you did before, answer the question seven times, choosing the portions of food that are your first, second, third, fourth, fifth, sixth and seventh best choices. Note: Think carefully, now that there are three openings, the situation is more complex.
@@
Upon completion of the cognitive test, each of the $N=213$ admitted participants distributed $\sum_{i=1}^{7}i=28$ points to $7$ different cells for each of the seven tasks. Thus, a total of $N_{\rm tot}=28N=5964$ datapoints were assigned to the $n=43$ cells of the detection screen. If $N(x|1)$, $N(x|2)$ and $N(x|3)$ denote the number of points assigned to cell $x$, when only slit-$1$, slit-$2$ and slit-$3$ are open, respectively, dividing them by $N_{\rm tot}$ one obtains the corresponding one-slit experimental probabilities $P(x|1)$, $P(x|2)$ and $P(x|3)$. One does the same with the numbers $N(x|12)$, $N(x|23)$ and $N(x|13)$, when only ‘slit-$1$ and $2$’, ‘slit-$2$ and $3$’, and ‘slit-$1$ and $3$’ are open, respectively, and with the number $N(x|123)$, when all three slits are open, obtaining in this way the two-slit experimental probabilities $P(x|12)$, $P(x|23)$ and $P(x|13)$, and the three-slit probability $P(x|123)$, $x=1,\dots,n$. All these probabilities are reported in Table 1, with the Sorkin parameter $\epsilon(x|123)$ and its normalized version $\kappa(x|123)=P_{\rm max}^{-1}\,\epsilon(x|123)$, $P_{\rm max}=\sup_{x}P(x|123)$.
Note that assuming a binomial distribution, the maximum value for the standard error is ${\rm SE}(19)\approx 0.00382$, where $x=19$ corresponds to the maximum probability $P_{\rm max}=P(19|123)\approx 0.0961$. If we consider a confidence level of $95\%$, the $z$-value is $1.96$ and the confidence interval becomes: $P(19|123)\approx 0.0961\pm 1.96\cdot 0.00382=0.0961\pm 0.0075$; see also Figure 5.
@@
Table 1: The seven experimental probabilities and the non-normalized and normalized Sorkin parameters $\epsilon(x|123)$ and $\kappa(x|123)=P_{\rm max}^{-1}\,\epsilon(x|123)$, $P_{\rm max}=\sup_{x}P(x|123)$, for the $43$ cells forming the detection screen.
@@ 5 Analysis of the experimental data
In this section, we analyze the data from our experiments and compare them with those typically obtained in physics laboratories. We begin by summarizing some of the values of the Sorkin parameter reported in the literature. Note that to allow for an easier comparison, normalized versions of the parameter are usually considered. In [32], a normalization factor $\delta(x|123)=|I(x|12)|+|I(x|23)|+|I(x|13)|$ was used. Performing a three-slit experiment with photons, $\epsilon^{\prime}(x|123)\equiv\delta^{-1}(x|123)\epsilon(x|123)$ was measured at the central maximum of the triple-slit interference pattern. The following values were obtained: $\epsilon^{\prime}(x|123)=0.0073\pm 0.0018$, when using a laser source and a power meter for detection; $\epsilon^{\prime}(x|123)=0.0034\pm 0.0038$, when attenuating the laser to single-photon level and a silicon avalanche photo-diode for detection; $\epsilon^{\prime}(x|123)=0.0064\pm 0.0119$, when using the heralded single-photon source. Hence, overall, they found $\epsilon^{\prime}(x|123)$ to be between $-0.0055$ and $0.0184$. In [34], an experiment with a three-path interferometer was performed and for the intensity maximum of the triple-slit interference pattern the following values were found: $\epsilon^{\prime}(x|123)=-0.0015\pm 0.0029$; $\epsilon^{\prime}(x|123)=-0.0050\pm 0.0018$; $\epsilon^{\prime}(x|123)=-0.0065\pm 0.00198$; $\epsilon^{\prime}(x|123)=-0.0142\pm 0.0010$, for increasing intensities at the three-path maximum. Hence, overall, the authors found $\epsilon^{\prime}(x|123)$ to be between $-0.0152$ and $0.0014$.
@@
In [40], it was proposed to use the intensity at the central maximum as a normalization factor: $\kappa(x|123)\equiv P_{\rm max}^{-1}\,\epsilon(x|123)$, $P_{\rm max}=\sup_{x}P(x|123)$. [^1] Note that from an algebraic point of view, $\epsilon(x|123)$ can range from $-2$ to $+2$. Since the minimal value of $P_{\rm max}$ is $1/n$, $\kappa(x|123)$ can range from $-2n$ to $+2n$. Using the Feynman path integral formalism in the thin-slit approximation, it was shown that $\kappa(x|123)$ is strongly dependent on certain experimental parameters and increases with an increase in wavelength. For the parameters used in [32], $\kappa(x|123)$ at the central maximum was found to be of the order of $10^{-6}$. However, for an incident beam of wavelength $4\,{\rm cm}$ (microwave regime), slit width of $120\,{\rm cm}$ and inter-slit distance of $400\,{\rm cm}$, their theoretical estimate of $\kappa(x|123)$ was of the order of $10^{-3}$. In [38], an upper limit for $|\kappa(123)|$ was derived, in the regime where the slit width $w$ and the inter-slit distance $d$ are such that $d\gg w$, and $w\gg{\lambda\over 2\pi}$, with $\lambda$ the wavelength: $|\kappa_{\rm max}(123)|\approx 0.03\lambda^{3\over 2}d^{-{1\over 2}}w^{-1}$.
@@
In [41], it was shown that looped trajectories of photons are due to the near-field component of the wavefunction, which leads to an interaction among the three slits. Hence, by modifying the dimensions of the three slit structure or by changing the wavelength of the optical excitation, the authors were able to increase the probability of occurrence of these trajectories, obtaining for instance values of $\kappa(x|123)$, for the central maximum, ranging from $-0.35$ to $0.15$. In their experiment they used a classical light from a tunable diode laser with wavelengths from $750\,{\rm nm}$ to $790\,{\rm nm}$, slit width $w=215\,{\rm nm}$, and inter-slit distance $d=4.6\,\mu{\rm m}$. In [43], experiments were performed in the microwaves domain, using pyramidal horn antennas as sources. A source-detector separation of $2.5\,{\rm m}$ was used ($1.25\,{\rm m}$ between source and slits plate and same distance between the slit plate and the detection screen). Non-zero values of $\kappa(x|123)$ of the order of $10^{-2}$ were measured, much above the error bound.
@@
Having summarized the experimental values reported in the physics literature, we can proceed to analyze our cognitive data, which we have represented graphically in Figures 6 and 9. Those for the three one-slit experiments, described in the left column of Figure 6, do not present any problems of interpretation. The question posed to the participants did not represent a situation of doubt, as there was certainty about the slit (the opening) from which the quantum entity (the animal) would emerge, and this explains the single peak forming in front of each of the slits.
@@
There is a further aspect that is particularly evident in the graph of $P(x|13)$, where the two slits are at a distance twice as far apart compared to the other two experiments. There are two additional secondary peaks in correspondence with the positions of the slits. These can also be glimpsed in the graph of $P(x|23)$. This choice doesn’t seem rational, but it was favored by a sufficient number of participants to make it emerge in the structure. The positions corresponding to the slits seem to be attractive in some way, although they tend to minimize rather than maximize doubt. Therefore, pure logical-rational thought processes were not the only ones involved in the decision-making process..
@@
$L_{1}=\sqrt{L^{2}+\left(x+d\right)^{2}},\quad L_{2}=\sqrt{L^{2}+x^{2}},\quad L_{3}=\sqrt{L^{2}+\left(x-d\right)^{2}}$ (9)
@@
To estimate the value of an effective wavelength $\lambda$ involved in the formation of the single-slit peaks, we have to use the graphs where only single answers are considered, otherwise the width of the peaks arbitrarily depends on the score assignment applied during analysis. We see in Figure 8 (fourth line) that peaks have a width of approximately $\Delta x\approx 3$. If we use the Fraunhofer diffraction formula (see Appendix A) $\Delta x\approx 2{\lambda L\over a}$, where $a$ is the width of the slit, since we have $a=1$ and $L=12$, the effective wavelength is $\lambda\approx a{\Delta x\over 2L}\approx{3\over 24}\approx 0.125$. On the other hand, If we apply the formula (see Appendix A) $\lambda\approx{xd\over L}$, giving the approximate distance $x$ at which a second maximum (the first being that of the central fringe) for the probability forms in the far field approximation, [^2] However, as we will discuss later in this article, this approximation does not adequately describe our experiment, given that the nullity of Sorkin’s parameter is strongly violated. considering (see Figure 6) $d=6$ and $x\approx{d\over 2}=3$, we find an effective wavelength $\lambda\approx{d^{2}\over 2L}=1.5$, which is clearly not compatible with the previously determined $\lambda\approx 0.125$ describing an effective diffraction mechanism. This means that there is no way to associate a single effective wavelength to our cognitive experiments, and this would be an important difference with the situation in physics laboratories (see also the discussion in Section 6).
@@
Figure 9: (color online) Right: Plot of the normalized Sorkin parameter $\kappa(x|123)$ (blue color). The probability $P(x|123)$, multiplied by a factor of $10$, is also plotted (red color), to give a clearer understanding of the modulation in the plot. A continuous approximation of the discrete functions was used to enhance interpretability in the visualization. Left: Plots of the normalized Sorkin parameter $\kappa(x|123)$ for different choices of the scores attributed to the participants’ choices (uniform, linear, exponential and single impact); see Figure 8).
We can observe that the values obtained are much higher than those typically obtained in physics laboratories. This reveals the presence of significant deviations of the probability $P(x|123)$ both from the classical situation and from the idealized quantum situation, when the principle of superposition can be applied naively. In other words, in their decision-making, participants were also guided by considerations that led to an overall assessment of the situation, taking into account the three slits together in their reasoning, rather than two at a time. This generated significant third-order interference-like contributions, which in physics can only be obtained in near-field situations, i.e., when the detection screen is not located far from the slits and/or the slits are very close together (distances on the order of the wavelength or less), so that non-radiative modes, such as evanescent waves, can easily mediate coupling between the slits and the detector, and/or between the slits.

### v3

@@ 4 The cognitive triple-slit experiment
In this section, we explain how our cognitive experiment was conducted in practice. For simplicity, we only considered a one-dimensional detection screen. Note that the typical interference fringes are a consequence of the slits’ geometry, in the sense that it is their height, and the fact that they are parallel to each other, that are responsible for the fringe appearance, while it is their distance from each other, and width, that are responsible for the alternation between high and low impact zones and diffraction phenomena. For the purposes of our experiment, however, it was not necessary to investigate the vertical dimension of the slits.
More precisely, our detection screen consisted of $n=8d+3$ elementary cells of unit length, where $d=5$ corresponds to the distance between the slits (the number of cells separating them) in the slit plate. The latter was placed at a distance of $L=12$ cells from the detection screen, which was formed of $n=43$ cells, with discrete positions $x=1,\dots,n$. The slits were chosen to be of unit width and placed in correspondence with cells with numbers $n=16,22,28$ (see Figures 5 and 7).
@@
More precisely, after agreeing to participate, each participant had to read an introductory text explaining the context of the experiment. Note that the introductory text, and subsequent explanatory texts, did not mention that it was a cognitive simulation of a quantum physics experiment, so as not to condition their choices on the basis of possible prior knowledge of the double-slit and possibly triple-slit experiments. For this reason, the different situations were presented to them in purely imaginative terms, as a description of an animal looking for food to eat (see below). In addition to the introductory text, each of the tasks was preceded by an explanation of the criterion by which they had to select a certain number of points (cells) on the detection screen. Basically, in each task participants had to select $7$ cells in sequence. The first selection corresponded to what was considered the best cell to answer the question asked, receiving (during analysis) a score of $7$ points. The second selected cell, which could not be the same because once a cell was selected the system disabled it and could not be selected a second time, received a score of $6$ points, and so on, until the seventh selected cell, which received only $1$ point. In other words, respondents answered the questions using a rank-ordering procedure, with scores from $7$ to $1$ assigned during analysis. This is of course not exactly what happens in a physics laboratory, where the measuring instrument records only one impact at a time. If each subject had been asked to select only one cell per task, however, the number of data collected might have been insufficient. Nevertheless, the patterns highlighted in this work remain identifiable regardless of the score that is attributed to the sequential choices of the participants, as we will explain in Section 5.
@@
For the single-slit experiments, participants were asked to select cells that were the best examples of impacts caused by particles emerging from the slit in question. For the experiments with two and three slits, they were asked to select cells that were the best examples of a situation of doubt about the slit from which the particle emerged. To best convey these situations, the various questions were posed in the form of a bet. The source of particles behind the slit plate was not described, as this would have complicated the understanding of the experiment and introduced considerations of the position of the source relative to the three slits. For the participants, each slit played the role of a potential source, with the constraint that only one at a time could emit a particle. More precisely, following some preliminary questions, participants had to read the following text (see Figure 5): This test evaluates how you interpret situations involving varying degrees of uncertainty. You will be presented with different situations and for each one you will be asked some questions. Although a question may seem similar to the previous one, you must always remain focused and answer as if it were the first time. The following description ensued: There is a large rectangular area. On one side there is a row of $43$ identical portions of food. On the opposite side is a wall with one or more openings. A hungry animal emerges from an opening and moves towards the line of food. Once at the line, the animal eats the portion of food in front of it and goes away. You have no information about the nature of the animal. You don’t know if the animal can sense the food, and if it can, you don’t know which sense it uses (sight, smell, etc.). This means that its path could even be exploratory or erratic.
@@
Participants had then to execute their first task, answering the following question: In this situation, the animal can only come out of one opening. You have bet a lot of money that you can guess which portion of food the animal will eat. What is your choice? To select a portion of food, click on it. Next, they were asked to provide their best second choice, and so on, until the seventh. Then, the same situation was presented to them two more times, changing the place of the opening. Following these three single-opening (single-slit) experiments, participants had to move to their fourth task, which was different from the previous ones.
The instruction for the fourth task was as follows: Many things have changed in this new situation: there are now two openings from which the animals can come out; now is you who decide which portions of food were eaten by the animals; each time a portion of food has been eaten, and the animal has left, another person, your opponent, will observe the situation and try to guess which opening the animal came from, taking into account which portion of food has been eaten; you have bet a lot of money that your opponent will not guess correctly; which portion of food should you choose for the animal to eat, so that your opponent cannot guess which opening it came from, and you win a lot of money? To select a portion of food, click on it. Note: Your opponent understands that there is uncertainty as to how the animal senses and reaches the food. In other words, in his/her reasoning, your opponent takes into account the fact that the animal may follow exploratory or irregular paths. [^3] Following Aerts’ conceptuality interpretation of the double-slit experiment (see Section 1), our experiment was designed to prompt participants to actualize outcomes in line with the idea that a detector selects the impact points most representative of the uncertainty inherent in the experimental context. This accounts for the apparent difference between the one-slit and double-slit situations, with an opponent introduced only in the latter. The opponent functioned merely as a stratagem to compel participants to focus on doubt, i.e., to select those cells that are most representative of the uncertainty about the opening from which the animal emerged. In the single-slit case, no opponent is needed, since the animal necessarily emerges from the only available opening. Again, participants had to provide their best seven choices in order, and the situation was then repeated twice with the two openings relocated.
Following these three double-opening (double-slit) experiments, participants had to move to their seventh and last task. The instruction for it was as follows: This is the last variation. There are now three openings for the animals to come out of. Which portions of food should you choose for the animals to eat, so that your opponents (different each time) cannot guess which opening they came from, and you win a lot of money? As you did before, answer the question seven times, choosing the portions of food that are your first, second, third, fourth, fifth, sixth and seventh best choices. Note: Think carefully, now that there are three openings, the situation is more complex.
@@
Upon completion of the cognitive test, each of the $N=213$ admitted participants distributed $\sum_{i=1}^{7}i=28$ points to $7$ different cells for each of the seven tasks. Thus, a total of $N_{\rm tot}=28N=5964$ datapoints were assigned to the $n=43$ cells of the detection screen. If $N(x|1)$, $N(x|2)$ and $N(x|3)$ denote the number of points assigned to cell $x$, when only slit-$1$, slit-22 and slit-$3$ are open, respectively, dividing them by $N_{\rm tot}$ one obtains the corresponding one-slit experimental probabilities $P(x|1)$, $P(x|2)$ and $P(x|3)$. One does the same with the numbers $N(x|12)$, $N(x|23)$ and $N(x|13)$, when only ‘slit-$1$ and $2$’, ‘slit-22 and $3$’, and ‘slit-$1$ and $3$’ are open, respectively, and with the number $N(x|123)$, when all three slits are open, obtaining in this way the two-slit experimental probabilities $P(x|12)$, $P(x|23)$ and $P(x|13)$, and the three-slit probability $P(x|123)$, $x=1,\dots,n$. All these probabilities are reported in Table 1, with the Sorkin parameter $\epsilon(x|123)$ and its normalized version $\kappa(x|123)=P_{\rm max}^{-1}\,\epsilon(x|123)$, $P_{\rm max}=\sup_{x}P(x|123)$. Note that assuming a binomial distribution, the maximum value for the standard error is ${\rm SE}(19)\approx 0.00382$, where $x=19$ corresponds to the maximum probability $P_{\rm max}=P(19|123)\approx 0.0961$. If we consider a confidence level of $95\%$, the $z$-value is $1.96$ and the confidence interval becomes: $P(19|123)\approx 0.0961\pm 1.96\cdot 0.00382=0.0961\pm 0.0075$; see also Figure 6.
@@
Table 1: The experimental probabilities for the $43$ cell and the Sorkin parameters $\epsilon(x|123)$ and $\kappa(x|123)=P_{\rm max}^{-1}\,\epsilon(x|123)$, $P_{\rm max}=\sup_{x}P(x|123)$.
@@ 5 Analysis of the experimental data
In this section, we analyze the data from our experiments and compare them with those typically obtained in physics laboratories. We begin by summarizing some of the values of the Sorkin parameter reported in the literature. Note that to allow for an easier comparison, normalized versions of the parameter are usually considered. In [33], a normalization factor $\delta(x|123)=|I(x|12)|+|I(x|23)|+|I(x|13)|$ was used. Performing a three-slit experiment with photons, $\epsilon^{\prime}(x|123)\equiv\delta^{-1}(x|123)\epsilon(x|123)$ was measured at the central maximum of the triple-slit interference pattern. The following values were obtained: $\epsilon^{\prime}(x|123)=0.0073\pm 0.0018$, when using a laser source and a power meter for detection; $\epsilon^{\prime}(x|123)=0.0034\pm 0.0038$, when attenuating the laser to single-photon level and a silicon avalanche photo-diode for detection; $\epsilon^{\prime}(x|123)=0.0064\pm 0.0119$, when using the heralded single-photon source. Hence, overall, they found $\epsilon^{\prime}(x|123)$ to be between 0.0055-0.0055 and $0.0184$. In [35], an experiment with a three-path interferometer was performed and for the intensity maximum of the triple-slit interference pattern the following values were found: $\epsilon^{\prime}(x|123)=-0.0015\pm 0.0029$; $\epsilon^{\prime}(x|123)=-0.0050\pm 0.0018$; $\epsilon^{\prime}(x|123)=-0.0065\pm 0.00198$; $\epsilon^{\prime}(x|123)=-0.0142\pm 0.0010$, for increasing intensities at the three-path maximum. Thus, overall, the authors found $\epsilon^{\prime}(x|123)$ to be between 0.0152-0.0152 and $0.0014$.
@@
In [41], it was proposed to use the intensity at the central maximum as a normalization factor: $\kappa(x|123)\equiv P_{\rm max}^{-1}\,\epsilon(x|123)$, $P_{\rm max}=\sup_{x}P(x|123)$. [^4] Note that $\epsilon(x|123)$ can range from 2-2 to +2+2. Since the minimal value of $P_{\rm max}$ is $1/n$, $\kappa(x|123)$ can range from 2n-2n to +2n+2n. These are of course purely algebraic bounds. Using the Feynman path integral formalism in the thin-slit approximation, it was shown that $\kappa(x|123)$ is strongly dependent on certain experimental parameters and increases with an increase in wavelength. For the parameters used in [33], $\kappa(x|123)$ at the central maximum was found to be of the order of $10^{-6}$. However, for an incident beam of wavelength $4\,{\rm cm}$ (microwave regime), slit width of $120\,{\rm cm}$ and inter-slit distance of $400\,{\rm cm}$, their theoretical estimate of $\kappa(x|123)$ was of the order of $10^{-3}$. In [39], an upper limit for $|\kappa(123)|$ was derived, in the regime where the slit width $w$ and the inter-slit distance $d$ are such that $d\gg w$, and $w\gg{\lambda\over 2\pi}$, with $\lambda$ the wavelength: $|\kappa_{\rm max}(123)|\approx 0.03\lambda^{3\over 2}d^{-{1\over 2}}w^{-1}$.
@@
In [42], it was shown that looped trajectories of photons are due to the near-field component of the wavefunction, which leads to an interaction among the three slits. Hence, by modifying the dimensions of the three slit structure or by changing the wavelength of the optical excitation, the authors were able to increase the probability of occurrence of these trajectories, obtaining for instance values of $\kappa(x|123)$, for the central maximum, ranging from 0.35-0.35 to $0.15$. In their experiment they used a classical light from a tunable diode laser with wavelengths from $750\,{\rm nm}$ to $790\,{\rm nm}$, slit width $w=215\,{\rm nm}$, and inter-slit distance $d=4.6\,\mu{\rm m}$. In [44], experiments were performed in the microwaves domain, using pyramidal horn antennas as sources. A source-detector separation of $2.5\,{\rm m}$ was used ($1.25\,{\rm m}$ between source and slits plate and same distance between the slit plate and the detection screen). Non-zero values of $\kappa(x|123)$ of the order of $10^{-2}$ were measured, much above the error bound.
@@
Having summarized the experimental values reported in the physics literature, we can proceed to analyze our cognitive data, which we have represented graphically in Figures 7 and 10. Those for the three one-slit experiments, described in the left column of Figure 7, do not present any problems of interpretation. The question posed to the participants did not represent a situation of doubt, as there was certainty about the slit (the opening) from which the quantum entity (the animal) would emerge, and this explains the single peak forming in front of each of the slits.
@@
There is another aspect that is particularly evident in the graph of $P(x|13)$, where one can observe two additional secondary peaks in correspondence with the positions of the slits. These can also be glimpsed in the graph of $P(x|23)$. A sufficient number of participants favored this choice to make it emerge in the structure. The positions corresponding to the slits appear to be attractive in some way. Is this a rational or irrational choice? There are two possible interpretations. One is to consider that this type of choice minimizes rather than maximizes doubt. This would mean that rational thought processes were not the only ones involved in the decision-making. The other interpretation is that, since the animal’s path could be exploratory or erratic, placing the food portion in correspondence with the opening (i.e., the slit) could be an equally valid way to deceive the opponent. Indeed, if we assume an erratic animal, then the reasoning that maximizes the distance between the food portion and the opening is no longer valid. In other words, strategies could consider doubt expressed not only at the level of the position of the food, but also at the level of its meaning in relation to the animal’s exploratory behavior.
@@
$L_{1}=\sqrt{L^{2}+\left(x+d\right)^{2}},\quad L_{2}=\sqrt{L^{2}+x^{2}},\quad L_{3}=\sqrt{L^{2}+\left(x-d\right)^{2}}$
@@
To estimate the value of an effective wavelength $\lambda$ involved in the formation of the single-slit peaks, we have to use the graphs where only single answers are considered, otherwise the width of the peaks arbitrarily depends on the score assignment applied during analysis. We see in Figure 9 (fourth line) that peaks have a width of approximately $\Delta x\approx 3$. If we use the Fraunhofer diffraction formula (see Appendix A) $\Delta x\approx 2{\lambda L\over a}$, where $a$ is the width of the slit, since we have $a=1$ and $L=12$, the effective wavelength is $\lambda\approx a{\Delta x\over 2L}\approx{3\over 24}\approx 0.125$. On the other hand, If we apply the formula $\lambda\approx{xd\over L}$ (see Appendix A), giving the approximate distance $x$ at which a second maximum (the first being that of the central fringe) for the probability forms in the far field approximation, [^5] Note however that, as we will discuss later in the article, this approximation does not adequately describe our experiment, given that the nullity of Sorkin’s parameter is strongly violated. considering $d=6$ and $x\approx{d\over 2}=3$ (see Figure 7), we find an effective wavelength $\lambda\approx{d^{2}\over 2L}=1.5$, which is clearly not compatible with the previously determined $\lambda\approx 0.125$ describing an effective diffraction mechanism. This means that there is no way to associate a single effective wavelength to our cognitive experiments, and this would be an important difference with the situation in physics laboratories (see also the discussion in Section 6).
@@
Figure 10: (color online) Top: Plot of the normalized Sorkin parameter $\kappa(x|123)$ (blue color). The probability $P(x|123)$, multiplied by a factor of $10$, is also plotted (red color), to give a clearer understanding of the modulation in the plot. A continuous approximation of the discrete functions was used to enhance interpretability in the visualization. Bottom: Plots of the normalized Sorkin parameter $\kappa(x|123)$ for different choices of the scores attributed to the participants’ choices (see Figure 9): uniform (light gray color), linear (blue color), exponential (magenta color) and single impact (pink color).
The values obtained are much higher than those typically obtained in physics laboratories. This reveals the presence of significant deviations of the probability $P(x|123)$ both from the classical situation and from the idealized quantum situation, when the principle of superposition can be applied naively. In other words, according to our quantum modeling of the experiment, in their decision-making participants were also guided by considerations that led to an overall assessment of the situation, taking into account the three slits together in their reasoning, rather than two at a time. This generated significant third-order interference-like contributions, which in physics can only be obtained in near-field situations, i.e., when the detection screen is not located far from the slits and/or the slits are very close together (distances on the order of the wavelength or less), so that non-radiative modes, such as evanescent waves, can easily mediate coupling between the slits and the detector, and/or between the slits.